OpenStudy (anonymous):

expand cos4x. I get that you get cos(2(2x)) but I don't know where to go from there.

zepdrix (zepdrix):

Is it \(\large \cos(4x)\) or \(\large \cos^4(x)\) ?

OpenStudy (anonymous):

the first one

zepdrix (zepdrix):

Hmm ok I guess we would have to apply the `Double Angle Formula for Cosine`. Remember this one? :)\[\large \cos (2x)=\cos^2x-\sin^2x\] We could also approach this using the `Angle Sum Formula for cosine`. \[\large \cos(a+b)=\cos a \cos b- \sin a \sin b\] Do either of these formula's look familiar?

OpenStudy (anonymous):

yes, I have learned both

zepdrix (zepdrix):

This step might not be totally necessary, but it'll simplify the problem hopefully. From the point you got to, let's make a quick substitution. \[\large \cos(2(2x))\]Let \(\large \color{royalblue}{2x=u}\) Which gives us, \(\large \cos(2\color{royalblue}{u})\) So we can apply the Double Angle Formula now right?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

So the formula is telling us, \[\large \cos(2\color{royalblue}{u})=\cos^2\color{royalblue}{u}-\sin^2\color{royalblue}{u}\]

zepdrix (zepdrix):

Let's go ahead and switch our u's back to x. That was just an intermediate step to help us through the process. \[\large =\cos^2(\color{royalblue}{2x})-\sin^2(\color{royalblue}{2x})\]

zepdrix (zepdrix):

Does the problem want us to get this down to an angle of x? :O If so you'll notice that we still have some work to do. We got our angle from 4x down to 2x by doing some work.

OpenStudy (anonymous):

yes, x

OpenStudy (anonymous):

how do you break that down more?

zepdrix (zepdrix):

Let's rewrite the squares in a different spot so this is easier to read.\[\large =(\color{green}{\cos 2x})^2-(\sin2x)^2\] So we didn't do anything yet, I just wrote the square on the outside so it's easier to see what's going on. Notice the green part. It looks like we can apply that same rule again! The Double Angle Formula for Cosine.

OpenStudy (anonymous):

ok I see.

zepdrix (zepdrix):

\[\large =(\color{green}{\cos^2x-\sin^2x})^2-(\sin2x)^2\] Understand what I did there? :)

OpenStudy (anonymous):

Oh, right. I do.

zepdrix (zepdrix):

So the green part is now a `binomial` with a square being applied to it. Understand how to expand the square out?

OpenStudy (anonymous):

so that would be \[(\cos ^{4}x -\sin ^{4}x )- (2 sinxcosx)^2 \]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

oh wait no

OpenStudy (anonymous):

go ahead

zepdrix (zepdrix):

Wooops, we can't just distribute the square to each term inside. We can't forget about the cross terms. When we see a square, what we actually have is this,\[\large =(\color{green}{\cos^2x-\sin^2x})(\color{green}{\cos^2x-\sin^2x})-(\sin2x)^2\] Sounds like maybe you caught your mistake though? :)

OpenStudy (anonymous):

that first part would be \[\cos^4 x -2\sin^2 x \cos^2 x -\sin ^4 x \]

zepdrix (zepdrix):

Yes very good! Oops, I think you have the sign on the sin^4 backwards. Negative times negative gives us a positive. And it looks like you remembered the `Double Angle Formula for Sine`.\[\large \cos^4 x -2\sin^2 x \cos^2 x +\sin ^4 x+(2\sin x \cos x)^2\]Hmm it looks like we can simplify things down a bit further from here.

OpenStudy (anonymous):

is the second part\[-(4\sin^2 x \cos^2 x) ?\]

zepdrix (zepdrix):

Yes good! Sorry I must've dropped the negative for some reason lol

zepdrix (zepdrix):

\[\large \cos^4 x\color{orangered}{-2\sin^2 x \cos^2 x} +\sin ^4 x\color{orangered}{-4\sin^2 x \cos^2 x}\]Hmm it looks like we can combine like terms, doesn't it?

OpenStudy (anonymous):

yes. let me write what I get for the finished product and you can tell me if I'm right.

OpenStudy (anonymous):

\[\cos^4 x -6 \sin^2 x \cos^2 x + \sin^4 x \]

zepdrix (zepdrix):

Yes, good job! :) It appears we were able to get all of our angles in terms of x, from 4x. So yay us!

OpenStudy (anonymous):

You were very helpful. Thank you!